A survey of 2,500 U.S. citizens yielded 1,600 who are in favor of gun control legislation. Estimate the proportion of all Americans who are in favor of gun control legislation using a 95% confidence interval.
Answer #1

Sample proportion = x/n = 1600/2500 = 0.64

For 95% confidence, z = 1.96

Hence,

95% confidence interval will be:

0.64 +1.96 + 0.64 +0.36 /1600

0.64 +0.0188

(0.6212, 0.6588)